OpenStudy (anonymous):

So how could I solve this equation? \[\frac{ 1 }{ 6-x } + \frac{ 2 }{ x+3 } = \frac{ 5x }{ x ^{2} - 3x - 18 }\] I know the answer is -15/4 but how could I solve it?

hero (hero):

1. First make sure the denominators are factored:

OpenStudy (anonymous):

but doesn't the \[x ^{2} - 3x - 18 = (x+3)(x-6)?\]

hero (hero):

\(\frac{ 1 }{ 6-x } + \frac{ 2 }{ x+3 } = \frac{ 5x }{ (x-6)(x+3) }\) Multiply both sides by (x-6)

hero (hero):

Let me know what you get

hero (hero):

I suppose you want me to post the full steps.

OpenStudy (anonymous):

ya i'm still figuring it out but it doesn't give what i want

hero (hero):

That's because you don't know how to implement the steps properly. I can show you on vyew.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

wait but how did you make (6-x) into (x-6)?

hero (hero):

Please don't get confused: 6-x = -(x-6)